OpenStudy (emogirlatemoooocow):

What is the reciprocal of 1|5? A. -5 B. - 1|5 C. 1|5 D. 5

OpenStudy (anonymous):

c

OpenStudy (maddy1251):

@ransondyara471 That isn't correct.

OpenStudy (nincompoop):

RECIPROCAL \(\huge \frac{p}{q} \rightarrow \color{red}{-} (\frac{q}{p}) \)

OpenStudy (maddy1251):

I am sure @nincompoop will have a good explanation for what the answer is.

OpenStudy (anonymous):

b

Nnesha (nnesha):

look at your previous post what you don't understand??

OpenStudy (anonymous):

they r always positive

OpenStudy (emogirlatemoooocow):

@Nnesha they didn't explain it well they basically gave out direct answer *_* .

OpenStudy (nincompoop):

\(\large -\frac{2}{3} \rightarrow \color{red}{-}(-\frac{3}{2}) = +\frac{3}{2}\)

Nnesha (nnesha):

there was an example reciprocal of a/b is b/a

OpenStudy (emogirlatemoooocow):

I'll just do the work on my own >.>

Nnesha (nnesha):

that's what you suppose to do. i hope you understand the exammple c/b reciprocal is b/c just flip the fraction that's it!